# Trustworthy Enhanced Multi-view Multi-modal Alzheimer’s Disease Prediction with Brain-wide Imaging Transcriptomics Data

### v1

@@ Abstract
Brain transcriptomics provides insights into the molecular mechanisms by which the brain coordinates its functions and processes. However, existing multimodal methods for predicting Alzheimer’s disease (AD) primarily rely on imaging and sometimes genetic data, often neglecting the transcriptomic basis of brain. Furthermore, while striving to integrate complementary information between modalities, most studies overlook the informativeness disparities between modalities. Here, we propose TMM, a trusted multiview multimodal graph attention framework for AD diagnosis, using extensive brain-wide transcriptomics and imaging data. First, we construct view-specific brain regional co-function networks (RRIs) from transcriptomics and multimodal radiomics data to incorporate interaction information from both biomolecular and imaging perspectives. Next, we apply graph attention (GAT) processing to each RRI network to produce graph embeddings and employ cross-modal attention to fuse transcriptomics-derived embedding with each imaging-derived embedding. Finally, a novel true-false-harmonized class probability (TFCP) strategy is designed to assess and adaptively adjust the prediction confidence of each modality for AD diagnosis. We evaluate TMM using the AHBA database with brain-wide transcriptomics data and the ADNI database with three imaging modalities (AV45-PET, FDG-PET, and VBM-MRI). The results demonstrate the superiority of our method in identifying AD, EMCI, and LMCI compared to state-of-the-arts. Code and data are available at https://github.com/Yaolab-fantastic/TMM.
@@
Brain transcriptomics, multimodal imaging, trustworthy learning, cross-modal attention, Alzheimer’s disease
@@ I Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that predominantly affects the elderly, characterized by the gradual deterioration of cognitive functions, such as memory, reasoning, and decision-making [1, 2]. Studies have demonstrated that the pathological changes associated with AD begin to manifest years or even decades before the earliest clinical symptoms are observed [3, 4].
Recent advances in acquiring biomedical data, coupled with the rapid advancements in machine learning and deep learning techniques, have significantly enhanced the development of computer-aided diagnosis of AD. The multifactorial nature of AD highlights the need for multi-modal learning, where numerous algorithms have been developed to capture and integrate complementary information from various modalities. Existing multimodal methods for AD prediction primarily focus on using brain imaging data [5, 6, 7, 8, 9]. For example, Qiu et al. [7] developed a local-aware convolution to capture intra-modality associations and employed a self-adaptive Transformer (SAT) to learn inter-modality global relationships. Zhang et al. [9] proposed an attention-based fusion framework to selectively extract features from MRI and PET branches.
@@
Although imaging modalities effectively quantify the structural and functional alterations in the brain, they do not account for the molecular mechanisms underlying these changes. A few studies have started to incorporate genotyping data to enhance model performance [10, 11, 12, 13]. For example, Zhou et al. [10] proposed an attentive deep canonical correlation analysis to integrate imaging modalities with genetic SNP data for diagnostic purposes. While genotypic data can characterize individuals, this upstream information does not directly explain cellular function and pathological states. Compared to genetics, downstream transcriptomics provides more direct insights by reflecting the actual activity and expression of genes within specific tissue contexts. This motivates us to incorporate brain-specific transcriptomics with imaging data into AD prediction.
@@
Besides improving representation learning, there has been a growing emphasis on developing various fusion strategies to capture the complementary information present across modalities [14, 15, 16, 17]. For example, Song et al. [16] developed functional and structural graphs using fMRI and DTI images, and combined the multimodal information into edges through an efficient calibration mechanism. Bi et al. [17] developed a community graph convolutional neural network to model brain region-gene interactions, using an affinity aggregation model to enhance interpretability and classification performance in AD diagnosis. Although effective, existing multimodal algorithms often lack a dynamical perception of the informativeness of each modality for different samples, which could otherwise enhance the trustworthiness, stability, and explainability of these methods. Evidence has shown that the informativeness of a modality typically varies across different samples [18, 19], which motivates efforts to model modality informativeness to enhance fusion effectiveness [20].
@@
To address the aforementioned issues, we propose a trustworthy enhanced multi-view multi-modal graph network (TMM) that integrates brain-wide transcriptomics knowledge and multi-modal radiomics data for predicting AD in an adaptive and trustworthy manner. In this context, multi-view refers to the incorporation of both transcriptomics and radiomics data, while multi-modal pertains to including different imaging modalities. As shown in Figure 1, we begin by constructing view-specific region of interest (ROI) co-function networks (RRIs) from brain-wide gene expression data and brain imaging data, to capture the molecular, structural, and functional relationships underlying different brain regions. Graph attention networks (GAT) are then applied to generate sample-wise embeddings, followed by cross-attention to fuse multi-view representations. In the fusion stage, we propose a novel true-false-harmonized class probability (TFCP) criterion to measure modality confidence, facilitating the adaptive perception of and response to variations in modality informativeness. Our main contributions can be summarized as follows:
@@
  • We propose a multi-view, multi-modal learning framework for AD prediction that integrates brain-specific transcriptomics and brain imaging data. This approach models the co-functional relationships of ROIs by combining insights from upstream molecular activities and downstream anatomical and functional characteristics.
@@
  • We propose a novel modality confidence learning strategy that estimates the harmonized true and false class probability to dynamically assess the modality informativeness. This design enables a more robust and trustworthy multimodal prediction.
@@
  • Our proposed model significantly outperforms state-of-the-art methods in predicting AD, LMCI, and EMCI. A series of ablation studies robustly validate the effectiveness of the proposed TMM framework. Furthermore, we identify important ROIs and demonstrate their functional roles in brain cognition.
@@ II Related work
Multimodal AD prediction. Existing multimodal methods for AD typically utilize brain imaging data (e.g., MRI, PET) because of their easy accessibility and early diagnosis potential. Common deep-learning architectures have been widely applied to learn embeddings of each individual modality, for example, convolutional neural network (CNN) [21], recurrent neural network (RNN) [22], autoencoder (AE) [23], generative adversarial network (GAN) [24] and transformer [25]. Most recently, graph neural networks (GNNs) and their variants have gained significant attention thanks to their natural advantages in capturing interaction patterns between ROIs and their adaptability to different data types [26].
@@
Besides imaging, genetics data have also been incorporated to enhance prediction performance [17, 13, 10]; however, genetic variants do not directly contribute to the pathogenesis of AD but instead influence the disease through downstream functional elements (e.g., gene expression). Although high-throughput sequencing techniques have greatly accelerated the acquisition of omics data, their application in brain disorders is challenging due to the invasive nature of sample collection. The scarcity of sample-level brain transcriptomics data inspires us to explore using valuable brain-wide transcriptomics data acquired from healthy individuals as prior knowledge for multimodal AD prediction.
@@
Confidence learning. Although various integration strategies have been proposed (e.g., early [27], intermediate [28], late [29], and hybrid [30] fusion), adapting to modal discrepancies remains challenging, particularly for heterogeneous and imbalanced multimodal data. By assessing and incorporating the predictive uncertainty of each modality into the fusion process, confidence learning has shown significant performance in multimodal prediction. Bayesian approaches introduce probability distributions over the model parameters to estimate uncertainty [31]. Ensemble methods train multiple models independently and use prediction variance as uncertainty [32]. Dropout-based methods, such as MC Dropout [33], leverage dropout regularization to obtain multiple predictions. Confidence calibration methods [34] aim to align the predicted probabilities with the true probabilities of correctness, thereby improving the reliability of uncertainty estimates. Han et al. [20] introduced the true class probability criterion to assess the uncertainty of each modality. Luo et al. [35] estimated the parameters of the Dirichlet distribution based on the prediction of each modality. Zheng et al. [36] assess the decision confidence of each modality by constraining the orthogonalities between each pair of predicted classes. In this paper, we designed a novel confidence calibration criterion to help estimate the informativeness of each modality for trustworthy integration.
@@
Figure 1: Framework overview. (a) Transcriptomics-specific RRI network (T-RRI) is constructed from the brain-wide gene expression data, and radiomics-specific RRI network is derived from each imaging modality. (b) Three imaging modalities, including VBM, FDG, and AV45, are employed in our analysis. (c) The TMM architecture: (i) For each sample, modality-specific T-RRI and R-RRI networks are constructed by mapping ROI measurements to each node. Multi-level GAT is applied to capture interactions within ROIs for each view, generating view-specific representations; followed by cross-view attention to create multi-view embeddings. (ii) TFCP-based confidence networks are designed to estimate the prediction confidence of each modality. (iii) Cross-modal attention mechanisms fuse multiple modalities to deliver the final predictive output. (d-f) illustrates the details of cross-view graph fusion, the TFCP mechanism, and the cross-attention mechanisms, respectively. (g) The identified ROI biomarkers are annotated for their functional connection and co-activation.
@@ III Methods
We now detail the proposed TMM framework. Initially, one transcriptomics-specific network (T-RRI) and $M$ radiomics-specific RRI networks (R-RRIs) are constructed from brain-wide gene expression data and each imaging modality dataset, respectively. For each imaging modality, ROI-level imaging measurements are assigned as node features to both the T-RRI and R-RRI networks, generating graph representations from transcriptomic and imaging views. GATs are then applied to each RRI network to generate view-specific embeddings. These embeddings are integrated using cross-modal attention mechanisms to form multi-view representations. Subsequently, a confidence learning network is specifically designed to enhance the reliability of the embeddings, which are further integrated using cross-modal attention for final prediction. The following sections provide further details on the construction of RRI networks, multi-view representation and fusion, trustworthy evaluation, and multi-modal integration. The overall architecture and important modules are illustrated in Figure 1.
III-A Sample-wise Co-functional RRI Network Construction
As shown in Figure 1(a), two types of RRI networks (T-RRI and R-RRI) are constructed from transcriptomics and imaging data, respectively, to model the co-functionality of ROIs and facilitate graph-based network implementation. Given the brain-wide gene expression matrix $\mathbf{T}\in\mathbb{R}^{n_{g}\times d}$, we derive the edge matrix $\mathbf{E}^{t}=\{e_{ij}^{t}\}\in\mathbb{R}^{d\times d}$ by thresholding the adjacency matrix, which measures the correlation coefficient for each pair of ROIs across all expressed genes:
$e_{ij}^{t}=\left\{\begin{array}[]{cl}1&\text{ if }r(\mathbf{t}_{i},\mathbf{t}_{j})\geq\lambda_{t}\\ 0&\text{ otherwise}\end{array}\right.,$ (1)
where $r(\mathbf{t}_{i},\mathbf{t}_{j})\!\!=\!\!\frac{\sum_{k=1}^{n_{g}}(\mathbf{t}_{ik}-\overline{\mathbf{t}_{i}})\cdot(\mathbf{t}_{jk}-\overline{\mathbf{t}_{j}})}{\sqrt{\sum_{k=1}^{n_{g}}(\mathbf{t}_{ik}-\overline{\mathbf{t}_{i}})^{2}}\cdot\sqrt{\sum_{k=1}^{n_{g}}(\mathbf{t}_{jk}-\overline{\mathbf{t}_{j}})^{2}}}$ computes the pearson correlation coefficient (PCC) of ROIs $\mathbf{t}_{i}$ and $\mathbf{t}_{j}$ expressions, and $\lambda_{t}$ is a hyperparameter for thresholding.
We apply the same strategy to each imaging modality, obtaining $M$ imaging-specific edge matrix $\mathbf{E}^{m}=\{e_{ij}^{m}\}\in\mathbb{R}^{d\times d}$ ($m\in[1,\dots,M]$), with each element $e^{m}_{i,j}$ representing the thresholded PCC (by $\lambda_{r}$) between imaging measures of each ROI pair across all $n$ samples. Across experiments, we set the thresholds $\lambda_{t}=0.2$ for transcriptomics-based networks and $\lambda_{r}=0.1$ for radiomics-based networks, respectively.
@@ III-B Multi-view Fusion for Modal-Specific Representation
View-specific representation. Given their ability to effectively aggregate neighboring relationships within graph data, we utilize GATs for multi-view sample representation learning. Specifically, GATs are applied to each T-RRI and R-RRI graph, enabling the sample representations from both transcriptomic and different imaging views. Illustrated with the $m$-th modality, we take $\mathcal{G}^{m}_{t,0}=(\mathbf{X}^{m},\mathbf{E}^{t})$ and $\mathcal{G}^{m}_{r,0}=(\mathbf{X}^{m},\mathbf{E}^{m})$ as input T-RRI and R-RRI graphs and stack graph attention layers with multi-head attention to build the GAT for each graph (as shown in Figure 1(d)). Each layer is defined as:
@@
We further adopted a multi-level strategy to enhance information aggregation between brain ROIs. Specifically, higher-level graphs $\mathcal{G}^{m}_{t,1}$ and $\mathcal{G}^{m}_{r,1}$ are generated by applying GAT to initial graphs $\mathcal{G}^{m}_{t,0}$ and $\mathcal{G}^{m}_{r,0}$. Then, we derived the next level graphs, $\mathcal{G}^{m}_{t,2}$ and $\mathcal{G}^{m}_{r,2}$ from $\mathcal{G}^{m}_{t,1}$ and $\mathcal{G}^{m}_{r,1}$ respectively. For each modality, graph embeddings from three levels are concatenated to produce view-specific representations $\mathbf{F}^{m}_{t}$ and $\mathbf{F}^{m}_{r}$.
Multi-view fusion. Within each modality, we introduce cross-view attention mechanisms to integrate representations from the T-RRI and R-RRI networks (as shown in Figure 1(d)), enabling them to complement and enhance each other:
@@
$\mathcal{L}_{\text{GAT}}^{m}=\sum\limits_{i=1}^{n}\mathcal{L}_{\text{CE}}(\hat{y}_{i}^{m},y_{i}),$ (4)
where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, ${n}$ is the number of training samples, $y_{i}$ is the true label, and $\hat{y_{i}^{m}}$ is the prediction from $\mathbf{Z}^{m}$.

### v2

@@ Abstract
Brain transcriptomics provides insights into the molecular mechanisms by which the brain coordinates its functions and processes. However, existing multimodal methods for predicting Alzheimer’s disease (AD) primarily rely on imaging and sometimes genetic data, often neglecting the transcriptomic basis of brain. Furthermore, while striving to integrate complementary information between modalities, most studies overlook the informativeness disparities between modalities. Here, we propose TMM, a trusted multiview multimodal graph attention framework for AD diagnosis, using extensive brain-wide transcriptomics and imaging data. First, we construct view-specific brain regional co-function networks (RRIs) from transcriptomics and multimodal radiomics data to incorporate interaction information from both biomolecular and imaging perspectives. Next, we apply graph attention (GAT) processing to each RRI network to produce graph embeddings and employ cross-modal attention to fuse transcriptomics-derived embedding with each imaging-derived embedding. Finally, a novel true-false-harmonized class probability (TFCP) strategy is designed to assess and adaptively adjust the prediction confidence of each modality for AD diagnosis. We evaluate TMM using the AHBA database with brain-wide transcriptomics data and the ADNI database with three imaging modalities (AV45-PET, FDG-PET, and VBM-MRI). The results demonstrate the superiority of our method in identifying AD, EMCI, and LMCI compared to state-of-the-arts. Code and data are available at https://github.com/Yaolab-fantastic/xx.
@@
Imaging transcriptomics, trustworthy learning, graph attention, cross-modal attention, Alzheimer’s disease
@@ I Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that predominantly affects the elderly, characterized by the gradual deterioration of cognitive functions, such as memory, reasoning, and decision-making [gaubert2021decision]. Studies have demonstrated that the pathological changes associated with AD begin to manifest years or even decades before the earliest clinical symptoms are observed [yao2019bioinfo].
@@
Recent advances in acquiring biomedical data, coupled with the rapid advancements in machine learning and deep learning techniques, have significantly enhanced the development of computer-aided diagnosis of AD. The multifactorial nature of AD highlights the need for multi-modal learning, where numerous algorithms have been developed to capture and integrate complementary information from various modalities. Existing multimodal methods for AD prediction primarily focus on using brain imaging data. https://doi.org/10.1016/j.media.2022.102698;DOI:10.1109/TMI.2024.3386937 Although imaging modalities effectively quantify the structural and functional alterations in the brain, they do not account for the molecular mechanisms underlying these changes. A few studies have also incorporated genotyping data into the prediction models to enhance their performance. ref: https://link.springer.com/chapter/10.1007/978-3-031-43895-0_64,https://doi.org/10.1016/j.media.2023.102883,10.1109/TMI.2022.3162870 For example, Bi et al. constructed brain region-gene networks using fMRI and SNP data and aggregated the multimodal information through a modified generative adversarial network [bi2022novel]. Venugopalan et al. [venugopalan2021multimodal] employed deep learning to integrally analyze MRI, SNP, and clinical test data, using stacked denoising auto-encoders and 3D-CNNs to enhance AD, MCI, and CN classification. While genotypic data can characterize individuals, this upstream information does not directly explain cellular function and pathological states. Compared to genetics, downstream transcriptomics provides more direct insights by reflecting the actual activity and expression of genes within specific tissue contexts. This motivates us to incorporate brain-specific transcriptomics with imaging data into AD prediction.
@@
Besides improving representation learning, there has been a growing emphasis on developing various fusion strategies to capture the complementary information present across modalities . For example, Song et al. [song2021graph] developed functional and structural graphs using fMRI and diffusion tensor imaging data, and combined the multimodal information into edges through an efficient calibration mechanism. Bi et al. [bi2023community] developed a community graph convolutional neural network to model brain region-gene interactions, using an affinity aggregation model to enhance interpretability and classification performance in AD diagnosis.Although effective, existing multimodal algorithms often lack a dynamical perception of the informativeness of each modality for different samples, which could otherwise enhance the trustworthiness, stability, and explainability of these methods. Evidence has shown that the informativeness of a modality typically varies across different samples [Dynamicsref[28, 49]], which motivates efforts to model modality informativeness to enhance fusion effectiveness .
@@
To address the aforementioned issues, we propose a trustworthy enhanced multi-view multi-modal graph network (TMM) that integrates brain-wide transcriptomics and multi-modal radiomics data for predicting AD in an adaptive and trustworthy manner. In this context, multi-view refers to the incorporation of both transcriptomics and radiomics data, while multi-modal pertains to including different imaging modalities. As illustrated in Figure LABEL:fig1_overview, we begin by constructing view-specific region of interest (ROI) co-function networks (RRIs) from brain-wide gene expression data and brain imaging data. These networks capture the molecular, structural, and functional relationships underlying different brain regions. Graph attention networks (GAT) are then applied to generate sample-wise embeddings, followed by cross-modal attention to fuse representations from different views. In the fusion stage, we propose a novel true-false-harmonized class probability (TFCP) mechanism to measure modal confidence, facilitating the adaptive perception of and response to variations in modality informativeness. Our main contributions can be summarized as follows:
@@
  • We propose a multi-view, multi-modal learning framework for AD prediction that integrates region-specific transcriptomics and brain imaging data. This approach models the co-functional relationships of ROIs by combining insights from upstream biomolecular activities and downstream anatomical and functional characteristics.
@@
  • We propose a novel modal confidence learning strategy that estimates the harmonized true and false class probability to dynamically assess the modality informativeness. This design enables a more robust and trustworthy multimodal prediction.
@@
  • Our proposed model significantly outperforms state-of-the-art methods in predicting AD and MCI. A series of ablation studies robustly validate the effectiveness of the proposed TMM framework. Furthermore, we identify important ROIs and demonstrate their functional roles in brain cognition.
@@ II Related work
Multimodal AD prediction. Existing multimodal methods for AD typically utilize brain imaging data (e.g., MRI, PET) because of their easy accessibility and early diagnosis potential. Common deep-learning architectures have been widely applied to learn embeddings of each individual modality, for example, convolutional neural network (CNN) , recurrent neural network (RNN) , autoencoder (AE) , generative adversarial network (GAN)  and transformer . Most recently, graph neural networks (GNNs) and their variants have gained significant attention thanks to their natural advantages in capturing interaction patterns between ROIs and their adaptability to different data types .
@@
Besides imaging, genetics data have also been incorporated to enhance prediction performance [e.g., Xia-anBi, LiShen&LifangHe]; however, genetic variants do not directly contribute to the pathogenesis of AD but instead influence the disease through downstream functional elements (e.g., gene expression). Although high-throughput sequencing techniques have greatly accelerated the acquisition of omics data, their application in brain disorders is challenging due to the invasive nature of sample collection. The scarcity of sample-level brain transcriptomics data inspires us to explore using brain-wide transcriptomics data from healthy individuals as prior knowledge for multimodal AD prediction.
@@
Confidence learning. Although various integration strategies have been proposed (e.g., early , intermediate , late , and hybrid  fusion), adapting to modal discrepancies remains challenging, particularly for heterogeneous and imbalanced multimodal data. By assessing and incorporating the predictive uncertainty of each modality into the fusion process, confidence learning has shown significant performance in multimodal prediction. Bayesian approaches introduce probability distributions over the model parameters to estimate uncertainty . Ensemble methods train multiple models independently and use prediction variance as uncertainty . Dropout-based methods, such as MC Dropout , leverage dropout regularization to obtain multiple predictions. Confidence calibration methods  aim to align the predicted probabilities with the true probabilities of correctness, thereby improving the reliability of uncertainty estimates. Han et al. [Han_2022_CVPR] introduced the true class probability criterion to assess the uncertainty of each modality. Luo et al. [luo2024teminet] estimated the parameters of the Dirichlet distribution based on the prediction of each modality. Zheng et al. [Zheng2023multilevel] assess the decision confidence of each modality by constraining the orthogonalities between each pair of predicted classes. In this paper, we designed a novel confidence calibration criterion to help estimate the informativeness of each modality for trustworthy integration.
@@
Figure 1: Framework overview. (a) Transcriptomics-specific RRI network (T-RRI) is constructed from the brain-wide gene expression data, and radiomics-specific RRI network is derived from each imaging data. (b) Three imaging modalities including VBM, FDG, and AV45 are employed in our analysis. (c) The TMM architecture. (i) For each sample, modality-specific T-RRI and R-RRI networks are constructed by mapping ROI measurements to each node. Multi-level GAT is applied to capture interactions within ROIs for each view, generating view-specific representations, followed by cross-view attention to create multi-view embeddings. (ii) The true-false-harmonized class probability (TFCP) is designed to estimate the prediction confidence for each modality. (iii) Cross-modal attention mechanisms fuse multiple modalities to deliver the final predictive output. (d-f) illustrates the details of cross-view graph fusion, the TFCP mechanism, and the cross-attention mechanisms, respectively. (g) The identified ROI biomarkers are annotated for their functional co-activation.
@@ III Methods
We now detail the proposed TMM framework. Initially, one transcriptomics-specific network (T-RRI) and $M$ radiomics-specific RRI networks (R-RRIs) are constructed from brain-wide gene expression data and each imaging modality dataset, respectively. For each imaging modality, ROI-level imaging measurements are assigned as node features to both the T-RRI and R-RRI networks, generating graph representations from transcriptomic and imaging views. Graph attention networks are then applied to each RRI network to generate view-specific embeddings, followed by cross-modal attention mechanisms to fuse multi-view representations. Subsequently, a confidence learning network is specifically designed to generate more reliable embeddings, with cross-modal attention employed to fuse these multimodal representations for final prediction. Below, we delve into the RRI co-functional networks construction, multi-view fusion, confidence learning module, and the multi-modal integration, detailed in Sections III-A,  III-B and  III-C, respectively. The overall architecture and important modules are illustrated in Figure 1.
III-A Sample-wise Co-functional Network Construction
Two types of RRI networks (T-RRI and R-RRI) are constructed from transcriptomics and imaging data, respectively, to model the co-functionality of ROIs and facilitate graph-based network implementation. Given the brain-wide gene expression matrix $\mathbf{T}\in\mathbb{R}^{n_{g}\times d}$, we derive the edge matrix $\mathbf{E}^{t}\in\mathbb{R}^{d\times d}$ by thresholding the adjacency matrix, which measures the correlation coefficient for each pair of ROIs across all expressed genes:
$\text{E}^{t}_{ij}=\left\{\begin{array}[]{cl}1&\text{ if }r(\mathbf{t}_{i},\mathbf{t}_{j})\geq\lambda\\ 0&\text{ otherwise}\end{array}\right.,$ (1)
where $r(\mathbf{t}_{i},\mathbf{t}_{j})=\frac{\sum_{k=1}^{g}(\mathbf{t}_{ik}-\overline{\mathbf{t}_{i}})\cdot(\mathbf{t}_{jk}-\overline{\mathbf{t}_{j}})}{\sqrt{\sum_{k=1}^{g}(\mathbf{t}_{ik}-\overline{\mathbf{t}_{i}})^{2}}\cdot\sqrt{\sum_{k=1}^{g}(\mathbf{t}_{jk}-\overline{\mathbf{t}_{j}})^{2}}}$ computes the Pearson correlation coefficient (PCC) between the expressions of ROIs $\mathbf{t}_{i}$ and $\mathbf{t}_{j}$, and $\lambda$ is a hyperparameter for thresholding.
We apply the same strategy to each imaging modality, obtaining $M$ imaging-specific edge matrix $\mathbf{E}^{m}\in\mathbb{R}^{d\times d}$ (where $m\in[1,\dots,M]$), with each element $\text{E}^{m}_{i,j}$ representing the PCC between imaging measures of each ROI pair across all $n$ samples. Across experiments, we set the thresholds $\lambda_{t}=0.2$ for transcriptomics-based networks and $\lambda_{r}=0.1$ for radiomics-based networks, respectively.
@@ III-B Multi-view Fusion for Modal-Specific Representation
View-specific representation. Given their ability to effectively aggregate neighboring relationships within graph data, we utilize GATs for multi-view sample representation learning. Specifically, multi-level GATs are applied to each T-RRI and R-RRI graph, enabling the sample representations from both transcriptomic and different imaging modal views. Illustrated with the $m$-th modality, we take $\mathcal{G}^{m}_{t,0}=(\mathbf{X}^{m},\mathbf{E}^{t})$ and $\mathcal{G}^{m}_{r,0}=(\mathbf{X}^{m},\mathbf{E}^{m})$ as input T-RRI and R-RRI graphs and stack graph attention layers with multi-head attention to build the GAT for each graph. Each layer is defined as:
@@
We further adopted a multi-level strategy to enhance the aggregation of information within brain circuits. Specifically, higher-level graphs $\mathcal{G}^{m}_{t,1}$ and $\mathcal{G}^{m}_{r,1}$ are generated by applying the GAT processes to initial graphs $\mathcal{G}^{m}_{t,0}$ and $\mathcal{G}^{m}_{r,0}$. Then, we derived the next level graphs, $\mathcal{G}^{m}_{t,2}$ and $\mathcal{G}^{m}_{r,2}$ from $\mathcal{G}^{m}_{t,1}$ and $\mathcal{G}^{m}_{r,1}$ respectively. For each modality, graph embeddings from all three levels are concatenated to produce view-specific representations, denoted as $\mathbf{F}^{m}_{t}$ and $\mathbf{F}^{m}_{r}$.
Multi-view fusion. Within each modality, we introduce cross-view attention mechanisms to integrate representations from the T-RRI and R-RRI networks, enabling them to complement and enhance each other:
@@
$\mathcal{L}_{\text{GAT}}^{m}=\sum\limits_{i=1}^{n}\mathcal{L}_{\text{CE}}(\hat{\mathbf{y}}_{i}^{m},\mathbf{y}_{i}),$ (4)
where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, ${n}$ is the number of training samples, $\mathbf{y}_{i}$ is the true label, and $\hat{\mathbf{y}}_{i}^{m}$ is the prediction from $\mathbf{Z}^{m}$.